A concentration cell was constructed by immersing two silver electrodes in 0.05 M and 

1M AgNO3 solutions. Give the cell representation, write the cell reactions and calculate 

the emf of the cell.



1
Expert's answer
2021-01-11T03:47:11-0500

Cell representation:

"Ag^0|Ag^+(0.05M)||Ag^+(1M)|Ag^0"

Cell reactions:

A(-) Ag0 - e- = Ag+

C(+) Ag+ + e+ =Ag0

Calculation of the emf of the cell:

EMF = "\frac{0.0592}{z}*log( \frac{C_2}{C_1})=\frac{0.0592}{1}*log( \frac{1}{0.05})=0.077V"